Can you please tell me if i got these correct?
Im doing a practice test to study for my real test. I would like to know if i got th questions correct or not.
Convert the linear equations to standard or slope-intercept form.
1.) 2x-3y=7 i got y=2/3x-7/3
2.) y=4/3x-10 i got 4x-3y=30
3.) .5x+y=12 i got y=.5x+12
4.) y=1/2x-2/3 i got 3x-6y=4
Solve the linear systems of equations using the elmination or addition/multiplication method.
5.) 9a+2b=10
9a-2b=8 I got a=1 b=1/2
6.) 3x+3y=15
4x-y=10 I got x=3 y=2
7.) 8x-2y=12
7x+3y=15 I got x=162/19 y=66/38
Solve the systems of linear equations using the substitution method.
8.) y=2x+3
2x+3y=10 I got (1/8,13/4)
9.) 5x-y=8
x+y=10 I got (1/2, 19/2)
10.) 12x+2y=7
y=x+5 i got (-3/14, 67/14)
11.) x=y+2
4y=3x-5 I got (-7, -9)
Solve the systems of linear equations using graphing method
12.) y=2x-13
y=-3x+7 I got (4, -5)
13.) 2x+y=10
x+3y=15 I got no solution, parallel lines
14.) 4x+2y=8
y= -2x+6 I got (3,2)
Word problem:
15.) Jill makes a blend of choclates for the bake sale. The dark chocolate is $6.00/ lb., and the milk chocolate is $4.00/ ib. Write an inequality shoing how Jill sells more than $25.00 for chocolate.
I got 6x+4y>25
If you can please tell me which ones I got wrong I would really apperciate it I really need to pass thi test and see if I need to strudy some more.
This would be a big help. Thanks

Found 2 solutions by Edwin McCravy, solver91311:
Answer by Edwin McCravy(20054)

These are wrong:

3.) .5x+y=12 i got y=.5x+12

7.) 8x-2y=12
7x+3y=15 I got x=162/19 y=66/38 

9.) 5x-y=8
x+y=10 I got (1/2, 19/2) 

11.) x=y+2
4y=3x-5 I got (-7, -9) 

13.) 2x+y=10
x+3y=15 I got no solution, parallel lines 

14.) 4x+2y=8
y= -2x+6 I got (3,2) 

Edwin

Answer by solver91311(24713)



Incorrect answers:

3. Sign error.

7. Not sure where you went wrong. Your value for is completly out to lunch and your value for is the value you should have gotten for . Try it again and send me the results in a reply. Hint: Multiply the first equation by 3 and the second one by 2 to eliminate the variable. Then, rather than trying to substitute that butt-ugly fraction back into one of the original equations, take the two original equations and multiply by -7 and 8 to eliminate the variable.

9. Not even close. Solve the second equation for



Then substitute into the first equation:



Solve for , etc. Pay attention to your signs

13. Dead wrong. Parallel lines have identical slopes. These two lines have different slopes so they must intersect.

15. I'm not sure what "...sells more than $25 for chocolate" means unless you are trying to say "...creates mixtures of lbs of dark chocolate and lbs of milk chocolate so that the value of the mixture is greater than $25." If that is the case, you did this one correctly.

Now, aside from going back and correcting all of the odd numbered problems that you missed, except of course for #15, I want you to put away the even-numbered problem answers you got from whatever source and then try the even numbered problems again.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism